Rochelle has 2.55 in dimes and quarters. She has eight5 more dimes than quarters.How many quarters does she have?
Answer by checkley71(8403)

THERE'S AN EXTRA 5 IN THE PROBLEM WHICH I WILL IGNORE.
25X+10(X+8)=255 OR 25X+10X+80=255 OR 35X+80=255 OR 35X=175 OR X=5
PROOF 25(5)+10(5+8)=255 OR 125+10(13)=255 OR 125+130=255